Miranda wants to give her 14 year old daughter $20000 when she turns 18. How much does she need to put in the bank now if the interest rate is 10 percent per year?
Answer by ankor@dixie-net.com(22740)

Miranda wants to give her 14 year old daughter $20000 when she turns 18.
How much does she need to put in the bank now if the interest rate is 10 percent per year?
:
Assuming the interest rate is compounded annually, Ao = initial amt
Ao*(1.1)^4 = 20000
1.4641*Ao = 20000
Ao = 20000/1.4641
Ao = $13,660.27